OpenStudy (anonymous):

The measure of two complimentary angles are in the ratio 1:4. What are the degree measures of the two angles.

OpenStudy (anonymous):

a. 18 deg. and 72 deg. b. 22.5 deg. and 67.5 deg. c. 36 deg. and 144 deg. d. 45 deg. and 135 deg.

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@M.D.Saxon

OpenStudy (anonymous):

What is the solution of the proportion 5/11=m/99? a. 1/45 b. 5/9 c. 9 d. 45

OpenStudy (paki):

@DesiFaith333 what are complimentary angles...? any idea...?

OpenStudy (anonymous):

um the same?

OpenStudy (paki):

Two angles are complementary if the sum of their angles equals 90 degrees... If one angle is known, its complementary angle can be found by subtracting the measure of its angle from 90 degrees.... GOT IT....

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so it would be d right?

OpenStudy (paki):

now here we have the ratios as 1:4... ok... so i am going to make an equation to solve the question... let suppose... Let the first angle be x Then second will be 4x then according to thhe definition of complementary angle.... we will have x+4x=90 ===> equation-1.... (agree with this)....

OpenStudy (anonymous):

yes

OpenStudy (paki):

hmmm... now solve equation-1 for "x"... what will we get....?

OpenStudy (anonymous):

The polygons are similar but not necessarily the drawn to scale. Find the value of x. a. 14 b. 21 c. 96 d. 504

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

is the first question answer d?

OpenStudy (paki):

for question What is the solution of the proportion 5/11=m/99?... it is 45...

OpenStudy (anonymous):

Thank you

OpenStudy (paki):

for question The measure of two complimentary angles are in the ratio 1:4. What are the degree measures of the two angles.... it is option A....

OpenStudy (anonymous):

Oh ok I was thinking D because of the 90 degree thing

OpenStudy (paki):

I WAS GUIDING YOU FOR ALL THE QUESTIONS, BUT OOPS YOU DIDN'T REPLIED ME PROPERLY.... :(

OpenStudy (anonymous):

oh im sorry :(

OpenStudy (anonymous):

Thank you though.

OpenStudy (paki):

pleasure...